OpenStudy (anonymous):

The formula r=2sqrt5L can be used to approximate the speed, r, in miles per hour, of a car that has left skid marks in the length, L, in feet. How far will a car skid at 80mph?

OpenStudy (turingtest):

\[r=2\sqrt{5L}\]is the equation?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

L=r^2/20=80^2/20=320 feet